# Addressing the Current Challenges of Quantum Machine Learning through Multi-Chip Ensembles

### v1

@@ Abstract
Quantum Machine Learning (QML) holds significant promise for solving computational challenges across diverse domains. However, its practical deployment is constrained by the limitations of noisy intermediate-scale quantum (NISQ) devices, including noise, limited scalability, and trainability issues in variational quantum circuits (VQCs). We introduce the multi-chip ensemble VQC framework, which partitions high-dimensional computations across smaller quantum chips to enhance scalability, trainability, and noise resilience. We show that this approach mitigates barren plateaus, reduces quantum error bias and variance, and maintains robust generalization through controlled entanglement. Designed to align with current and emerging quantum hardware, the framework demonstrates strong potential for enabling scalable QML on near-term devices, as validated by experiments on standard benchmark datasets (MNIST, FashionMNIST, CIFAR-10) and real world dataset (PhysioNet EEG).
@@ 1 Introduction
The past decade has witnessed significant advancements in quantum computing hardware, algorithms, and applications [1, 2, 3, 4, 5]. Among these, Quantum Machine Learning (QML) has emerged as a promising approach to harness quantum computing for real-world challenges, leveraging properties such as superposition and entanglement to achieve computational advantages over classical methods [6, 7]. Applications range from quantum simulation in chemistry [8, 9] and materials science [4, 10] to healthcare [11, 12], sensing [13, 14, 15], and high-energy physics [16, 17].
Despite these prospects, QML faces critical challenges due to the limitations of noisy intermediate-scale quantum (NISQ) devices [18, 19]. These devices, characterized by noise, limited coherence, and sparse qubit connectivity, pose significant barriers to the scalability and trainability of QML models, particularly variational quantum circuits (VQCs) [20]. The emergence of barren plateaus [21], where gradients vanish in the optimization landscape, further exacerbates these challenges.
To overcome these NISQ-era barriers, we introduce multi-chip ensemble VQCs—a modular architecture that splits the input across several small quantum chips and aggregates their measurements classically. This partitioning scales to high-dimensional data without enlarging any single chip; tames barren plateaus by capping inter-chip entanglement; boosts generalization through implicit regularization; dampens hardware noise because each subcircuit is shorter and errors average out. Instead of dissecting every possible design choice (e.g., circuit depth, quantum gate set, partition heuristics), we demonstrate theoretically and empirically that these advantages stem from the framework’s ability to control quantum entanglement. While direct experiments on real quantum hardware was infeasible, we validated the multi-chip ensemble framework under realistic depolarizing and amplitude-damping noise, mirroring today’s NISQ conditions. The scheme is fully compatible with existing hardware and fits the multi-chip roadmaps of IBM, IonQ, and Rigetti [22, 23, 24].
@@ 2.2 Limitations of VQC
VQCs are hybrid quantum-classical algorithms designed to utilize the computational power of NISQ devices. They parameterize quantum circuits to solve optimization problems, where a classical optimizer iteratively updates parameters to minimize a cost function. By leveraging quantum hardware for state preparation and measurement, and classical resources for optimization, VQCs bridge the gap between NISQ limitations and real-world computational demands. This hybrid approach has enabled applications in quantum chemistry, combinatorial optimization, and quantum machine learning, making VQCs a cornerstone of near-term quantum computing research [26].
@@
Barren plateaus—regions in the optimization landscape where gradients vanish—are a significant obstacle to VQC optimization [21, 20, 31]. Gradients often vanish exponentially with the number of qubits, making optimization difficult for many problems [32]. Noise and random parameter initialization exacerbate these challenges, creating uninformative and noisy loss landscapes that hinder effective parameter updates [33]. These factors reduce the generalizability of VQCs on complex datasets and limit their utility in quantum machine learning.
@@
Noise, inherent to NISQ hardware, affects all stages of VQC execution, from state preparation to measurements [26]. Noise arises from interactions between qubits and their environment, introducing errors in quantum gates, measurements, and state preparation. This limits circuit depth and computational accuracy [19]. Accumulated noise reduces circuit fidelity exponentially with depth, limiting the expressivity of VQCs and exacerbating barren plateau issues [33]. These combined effects restrict VQCs to shallow circuits, narrowing the range of problems they can address.
@@ 2.3.2 Limitations of existing Distributed QML strategies
Current distributed QML approaches face four principal limitations. First, circuit cutting techniques [44, 46] fragment large circuits into smaller segments but incur exponential sampling overhead, undermining their practical scalability. Second, communication-based methods [47] that integrate mid-circuit classical information exchange suffer from latency and coherence disruptions. Third, feature-based partitioning approaches [45] divide input data across multiple circuits with classical output aggregation, but lack theoretical foundations regarding trainability, generalization, and noise resilience. Fourth, quantum federated learning methods [48], while promising for privacy preservation, remain primarily heuristic without rigorous analysis of fundamental QML challenges.
@@ 2.3.3 Novel contribution of the present work
: The proposed framework processes high-dimensional data using ensembles of shallow circuits, enabling scalability without increasing qubit count per chip. Our architecture aligns with modular quantum hardware roadmaps (e.g., IBM, IonQ, Rigetti), making it forward-compatible with near-term devices [22, 23, 24].
@@ 3 Multi-Chip Ensemble VQCs
Figure 1: Comparison between Single-chip vs Multi-chip Ensemble VQCs
@@ 3.3 Entanglement in Multi-Chip Ensembles
Quantum entanglement—the non-classical correlation between quantum subsystems—fundamentally shapes the computational capabilities and limitations of QML models. When unitary operators $U(\boldsymbol{\theta})$ act on encoded quantum states $\rho(\boldsymbol{x})$, the resulting state $U(\boldsymbol{\theta})\rho(\boldsymbol{x})U^{\dagger}(\boldsymbol{\theta})$ can exhibit entanglement patterns that fundamentally determine the model’s expressibility, trainability, and generalizability.
A key feature of our multi-chip ensemble architecture is its controlled entanglement structure: quantum connections exist only among qubits within each chip, with no entanglement between qubits across different chips. This design choice leads to quantifiably lower global entanglement compared to equivalent single-chip circuits (see Appendix B for formal proofs).
While this reduction in entanglement restricts the accessible portion of the Hilbert space [52, 53, 54], it simultaneously addresses two critical challenges in quantum machine learning. First, it mitigates barren plateaus—regions where gradients vanish exponentially with system size [55, 56, 57]—by preventing the global entanglement patterns that trigger this phenomenon. Second, it reduces overfitting by limiting the model’s capacity to represent excessively complex functions [58], effectively serving as an implicit regularization mechanism.
Our approach demonstrates a fundamental principle: optimal quantum model performance requires calibrating entanglement appropriately rather than maximizing it. The multi-chip ensemble architecture provides a systematic framework for achieving this balance, as confirmed by our experimental results across diverse datasets.
@@ 4.1 Improved Scalability
The scalability of multi-chip ensemble VQCs offers significant advantages over conventional single-chip VQCs, particularly for processing high-dimensional data in machine learning applications. Single-chip VQCs face scalability limitations, as variational encoding requires $n$ qubits to process $n$-dimensional data. This often necessitates classical dimension reduction techniques, such as principal component analysis, autoencoders, or learnable neural networks, which can introduce information loss and degrade model performance (Figure 3(a)). While amplitude encoding theoretically enables encoding $2^{n}$-dimensional data with $n$ qubits by leveraging quantum superposition, its practical use is limited by deep circuit requirements and complex state preparation, making it infeasible for current NISQ hardware [27, 28, 29].
In contrast, multi-chip ensemble VQCs distribute the computational load across $k$ independent quantum chips, each processing $l=n/k$ dimensions of the input data (Figure 3(c)). This architecture eliminates the need for mandatory dimension reduction by scaling horizontally through additional chips rather than requiring larger single chips. For instance, a small quantum circuit with $10$ physical qubits can be combined $200$ times to conceptually create an ensemble model with $2000=10\times 200$ qubits, enabling $2000$-dimensional data processing using only $10$ physical qubits. Additionally, each subcircuit can specialize in specific feature subspaces, capturing nuanced patterns that might be lost with global dimension reduction.
The multi-chip design also enables advanced parallel data processing strategies. Correlated features can be grouped and processed on the same chip, while independent features are distributed across different chips. This natural partitioning of the feature space aligns well with many real-world datasets, where features often exhibit clustered correlations.
@@ 4.3 Improved Generalizability
Generalization—a model’s ability to perform well on unseen data—represents a critical challenge in machine learning that manifests uniquely in quantum systems. Our multi-chip ensemble approach provides a principled framework for optimizing generalization through controlled entanglement.
Quantum Bias-Variance Trade-off
In classical machine learning, the bias-variance trade-off balances model complexity against overfitting risk [67]. This fundamental principle extends to quantum systems, where increased circuit complexity reduces bias (underfitting) but potentially increases variance (overfitting) [68, 69]. Quantum models with appropriately calibrated model complexity can achieve exceptional generalization, even with limited training samples [70, 71], but excessive complexity leads to deteriorating test performance. This relationship creates a critical need for methods that can navigate this trade-off effectively in quantum circuits.
Entanglement as a Complexity Regulator
Our theoretical analysis, detailed in Appendix E, establishes that quantum entanglement directly modulates this bias-variance trade-off. We demonstrate that higher entanglement levels ($\gamma_{k}$) expand the representable function class, reducing bias in the learning process. However, this expanded expressibility simultaneously increases the risk of overfitting, captured by a complexity penalty $\Omega(\gamma_{k})$ in our generalization bounds. The optimal generalization performance occurs at an intermediate entanglement level that balances these competing factors. This relationship is formalized as:
@@
This formulation demonstrates that entanglement serves as a fundamental regulator of the quantum bias-variance trade-off, with important implications for quantum model design.
Multi-Chip Ensemble as Implicit Regularization
Our multi-chip ensemble architecture implements this theoretical insight by confining entanglement within chip boundaries. By partitioning $n$ qubits into $k$ independent chips, we reduce global entanglement levels proportionally to the number of chips ($\gamma_{k}\propto 1/k$). This design creates an implicit regularization mechanism: each chip maintains sufficient internal expressibility to capture relevant data patterns, while the absence of cross-chip entanglement prevents the variance surge associated with global entanglement. The architecture naturally positions the model near the optimal point on the bias-variance curve without requiring additional techniques like entanglement dropout [58].
@@ 4.4 Improved Noise Resilience
Our multi-chip ensemble framework inherently reduces both bias and variance of quantum errors simultaneously without additional resources. For a single-chip circuit with $n$-qubits and $N_{g}$ noisy gates, depolarizing-channel analysis shows that bias grows exponentially as $\exp(nN_{g}\varepsilon)$, while variance scales inversely with circuit runs as $1/N_{\text{cir}}$. When partitioning this workload across $k$ independent $l(=n/k)$-qubit chips, each chip contributes a significantly smaller bias term $\exp\!\bigl{(}\tfrac{n}{k}N_{g}\varepsilon\bigr{)}$. Summing across chips yields a total bias of $k\exp\bigl{(}\tfrac{n}{k}N_{g}\varepsilon\bigr{)}$, which remains exponentially smaller than single-chip bias for any $k>1$. Furthermore, because chip noise patterns are uncorrelated, their classical averaging reduces variance to $1/(kN_{\text{cir}})$.
@@ 5 Experiments
Figure 2: Experimental Results on MNIST. DimReduc 2-Chip and 4-Chip indicate multi-chip ensemble VQC models with classical dimension reduction. Ensemble 98-Chip denotes multi-chip ensemble VQC model without classical dimension reduction.
@@ 5.3 Trainability
Using the entangling capability measure [52], we quantified the quantum entanglement levels in single-chip and multi-chip ensemble VQCs with classical dimension reduction. For a total of 8 qubits, increasing the number of chips (i.e., dividing a large VQC into smaller subcircuits) reduced the overall entanglement while increasing the variance of gradients (Table LABEL:Table_Entanglement).
Since barren plateaus are characterized by exponentially small gradient variance, the higher gradient variance observed in multi-chip ensemble VQCs suggests a reduced risk of barren plateaus. These results, consistent with our theoretical analysis in Appendix C, demonstrate that the multi-chip ensemble approach can mitigate the risk of barren plateaus, thereby enhancing trainability.
@@ 6 Discussion
The multi-chip ensemble VQC framework effectively addresses key limitations of conventional single-chip quantum machine learning models. By distributing computations across smaller, independent quantum chips and leveraging classical postprocessing, this approach overcomes scalability constraints, reduces noise accumulation, and enhances both generalizability and noise resilience. Experimental results demonstrate that multi-chip ensembles outperform single-chip VQCs in processing high-dimensional data, even without classical dimension reduction, and achieve superior performance, lower generalization loss, and reduced quantum errors.
Multi-chip ensemble VQCs’ ability to directly process high-dimensional data without dimension reduction preserves data fidelity. Machine learning tasks in areas such as computer vision, natural language processing, and genomics frequently involve high-dimensional data, where complex feature interactions are critical for model performance. By maintaining these relationships, multi-chip ensemble VQCs can capture more sophisticated patterns compared to single-chip VQCs, which often sacrifice information to meet hardware limitations.
Future research on multi-chip ensemble VQCs could explore the design of heterogeneous subcircuits for individual chips. Tailoring subcircuit architectures or assigning specific feature subspaces to different chips may further optimize performance and enhance the framework’s adaptability to diverse tasks. Additionally, investigating advanced strategies for inter-chip communication and integration with emerging multi-chip quantum hardware will be critical for scaling the approach to even larger datasets and more complex problems.
The multi-chip ensemble VQC represents a forward-compatible and practical solution for advancing quantum machine learning, paving the way for robust, scalable applications in the NISQ era and beyond.

### v2

@@ Abstract
Practical Quantum Machine Learning (QML) is challenged by noise, limited scalability, and poor trainability in Variational Quantum Circuits (VQCs) on current hardware. We propose a multi-chip ensemble VQC framework that systematically overcomes these hurdles. By partitioning high-dimensional computations across ensembles of smaller, independently operating quantum chips and leveraging controlled inter-chip entanglement boundaries, our approach demonstrably mitigates barren plateaus, enhances generalization, and uniquely reduces both quantum error bias and variance simultaneously without additional mitigation overhead. This allows for robust processing of large-scale data, as validated on standard benchmarks (MNIST, FashionMNIST, CIFAR-10) and a real-world PhysioNet EEG dataset, aligning with emerging modular quantum hardware and paving the way for more scalable QML.
@@ 1 Introduction
Quantum Machine Learning (QML) has emerged as a promising approach to leverage quantum properties like superposition and entanglement for computational advantage in diverse domains [1, 2], from chemistry [3, 4] and materials science [5, 6] to healthcare [7, 8] and high-energy physics [9, 10]. However, QML faces critical challenges due to noisy intermediate-scale quantum (NISQ) hardware limitations [11, 12], including noise, limited coherence, sparse connectivity, and the emergence of barren plateaus—regions where gradients vanish exponentially with system size [13, 14]. These factors severely hamper scalability and trainability of QML models, particularly variational quantum circuits (VQCs) [15].
We introduce multi-chip ensemble VQCs to address these challenges through a novel modular architecture that processes input subsets on multiple independent quantum processors with classical aggregation of quantum measurement outputs. This approach offers four key advantages over previous distributed quantum strategies: (1) high-dimensional data processing without lossy dimension reduction; (2) barren plateau mitigation through controlled entanglement; (3) enhanced generalization via implicit regularization; and (4) simultaneous reduction of both bias and variance in quantum errors—an improvement over traditional error mitigation techniques that typically trade one for the other.
Our theoretical framework formally establishes these benefits through analysis of entanglement properties, gradient variance scaling, and noise propagation. Experimental validation demonstrates significant performance improvements over both classical baselines and single-chip QML approaches under realistic noise conditions. The architecture maintains compatibility with current hardware and aligns with modular quantum computing roadmaps [16, 17, 18], providing a practical pathway toward scalable QML on near-term quantum devices.
@@ 2.2 Limitations of VQC
VQCs are hybrid quantum-classical algorithms designed to utilize the computational power of NISQ devices. They parameterize quantum circuits to solve optimization problems, where a classical optimizer iteratively updates parameters to minimize a cost function. By leveraging quantum hardware for state preparation and measurement, and classical resources for optimization, VQCs bridge the gap between NISQ limitations and real-world computational demands. This hybrid approach has enabled applications in quantum chemistry, combinatorial optimization, and QML, making VQCs a cornerstone of near-term quantum computing research [20].
@@
Barren plateaus—regions in the optimization landscape where gradients vanish—are a significant obstacle to VQC optimization [13, 15, 14]. Gradients often vanish exponentially with the number of qubits, making optimization difficult for many problems [25]. Noise and random parameter initialization exacerbate these challenges, creating uninformative and noisy loss landscapes that hinder effective parameter updates [26]. These factors reduce the generalizability of VQCs on complex datasets and limit their utility in QML.
@@
Noise, inherent to NISQ hardware, affects all stages of VQC execution, from state preparation to measurements [20]. Noise arises from interactions between qubits and their environment, introducing errors in quantum gates, measurements, and state preparation. This limits circuit depth and computational accuracy [12]. Accumulated noise reduces circuit fidelity exponentially with depth, limiting the expressibility of VQCs and exacerbating barren plateau issues [26]. These combined effects restrict VQCs to shallow circuits, narrowing the range of problems they can address.
@@ 2.3.2 Limitations of existing Distributed QML strategies
Current distributed QML approaches face four principal limitations. First, circuit cutting techniques [37, 39] fragment large circuits into smaller segments but incur exponential sampling overhead, undermining their practical scalability. Second, communication-based methods [40] that integrate mid-circuit classical information exchange suffer from latency and coherence disruptions. Third, feature-based partitioning approaches [38] divide input data across multiple circuits with classical output aggregation, but restrict aggregation to a simple arithmetic mean and lack theoretical foundations regarding trainability, generalization, and noise resilience. Fourth, quantum federated learning methods [41], while promising for privacy preservation, remain primarily heuristic without rigorous analysis of fundamental QML challenges.
@@ 2.3.3 Novel contribution of the present work
We extend beyond simple summation aggregation of individual quantum circuit outputs [38] to general classical post-processing functions that can be optimized for specific tasks. The proposed framework processes high-dimensional data using ensembles of shallow circuits, enabling scalability without increasing qubit count per chip. Our architecture aligns with modular quantum hardware roadmaps (e.g., IBM, IonQ, Rigetti), making it forward-compatible with near-term devices [16, 17, 18].
@@ 3 Multi-Chip Ensemble VQCs
Figure 1: Comparison between Single-chip vs Multi-chip Ensemble VQCs. (a) Conventional single-chip approach processes entire input $\boldsymbol{X}$ through a single large VQC. (b) Multi-chip ensemble partitions input into subvectors ($\boldsymbol{x}_{1},...\boldsymbol{x}_{k}$), each processed by independent smaller VQCs on separate chips with outputs classically combined to produce $\boldsymbol{Y}$. This distributed architecture enhances scalability, trainability, and noise resilience.
@@ 3.3 Entanglement in Multi-Chip Ensembles
Quantum entanglement fundamentally determines QML models’ expressibility, trainability, and generalizability. Our multi-chip ensemble architecture introduces a controlled entanglement structure: quantum connections exist only within each chip, not between chips. This design creates quantifiably lower global entanglement compared to single-chip circuits (see Appendix B for formal proofs).
This controlled entanglement simultaneously addresses two critical QML challenges while accepting a constrained Hilbert space [46, 47, 48]. First, it mitigates barren plateaus [49, 50, 51] by preventing the global entanglement patterns that trigger this phenomenon. Second, it reduces overfitting by constraining the model’s representational capacity [52], effectively implementing implicit regularization.
Our approach demonstrates that optimal QML performance requires calibrating entanglement appropriately rather than maximizing it—a principle systematically implemented through our multi-chip ensemble architecture.
@@ 4.1 Improved Scalability
Single-chip VQCs face inherent scalability constraints: processing $n$-dimensional data typically requires $n$ qubits, often necessitating classical dimension reduction techniques that introduce information loss (Figure 3(a)). While amplitude encoding theoretically enables encoding $2^{n}$-dimensional data with $n$ qubits, its practical implementation requires prohibitively deep circuits and complex state preparation for current NISQ hardware [21, 22, 23].
Our multi-chip ensemble VQCs overcome this limitation by distributing computational load across $k$ independent quantum chips, each processing $l=n/k$ dimensions of the input data (Figure 3(c)). This horizontal scaling approach eliminates mandatory dimension reduction—a small circuit with 10 physical qubits, replicated 200 times, can process 2,000-dimensional data that would otherwise require 2,000 qubits in a single-chip implementation.
This architecture offers additional advantages through strategic feature allocation: correlated features can be grouped on the same chip while independent features are distributed across different chips, aligning naturally with the clustered correlation structure often present in real-world datasets. Each subcircuit can thus specialize in specific feature subspaces, preserving nuanced patterns that might be lost with global dimension reduction.
@@ 4.3 Improved Generalizability
Our multi-chip ensemble approach provides a principled framework for optimizing generalization through controlled entanglement, addressing the quantum manifestation of the classical bias-variance trade-off.
In quantum systems, increased circuit complexity reduces bias but potentially increases variance [61, 62]. Our theoretical analysis (Appendix E) establishes that quantum entanglement directly modulates this trade-off: higher entanglement levels ($\gamma_{k}$) expand the representable function class (reducing bias) while increasing overfitting risk (increasing variance), captured in our generalization bound:
@@
Multi-chip ensemble architecture implements this insight by confining entanglement within chip boundaries. By partitioning $n$ qubits into $k$ independent chips, we reduce global entanglement proportionally ($\gamma_{k}\propto 1/k$), creating an implicit regularization mechanism. Each chip maintains sufficient internal expressibility for pattern capture, while the absence of cross-chip entanglement prevents the variance surge associated with global entanglement, naturally positioning the model near the optimal point on the bias-variance curve without requiring additional techniques like entanglement dropout [52].
@@ 4.4 Improved Noise Resilience
Our multi-chip ensemble framework inherently reduces both bias and variance of quantum errors simultaneously without additional resources. For a single-chip circuit with $n$-qubits and $N_{g}$ noisy gates, bias grows exponentially as $\exp(nN_{g}\varepsilon)$, while variance scales inversely with circuit runs as $1/N_{\text{cir}}$. When partitioning this workload across $k$ independent $l(=n/k)$-qubit chips, each chip contributes a significantly smaller bias term $\exp\!\bigl{(}\tfrac{n}{k}N_{g}\varepsilon\bigr{)}$. Summing across chips yields a total bias of $k\exp\bigl{(}\tfrac{n}{k}N_{g}\varepsilon\bigr{)}$, which remains exponentially smaller than single-chip bias for any $k>1$. Furthermore, because chip noise patterns are uncorrelated, their classical averaging reduces variance to $1/(kN_{\text{cir}})$.
@@ 5 Experiments
Figure 2: Experimental Results on MNIST. (a) Training and validation loss across epochs, showing multi-chip ensembles outperform single-chip VQCs and classical baselines. (b) Generalization error, with lower values indicating better generalization. (c) Quantum error under simulated noise, demonstrating multi-chip models’ superior noise resilience compared to both standard and ZNE-mitigated single-chip VQCs. Ensemble 98-Chip processes full-dimensional data without classical reduction, while DimReduc models use multi-chip ensembles with dimension reduction.
(a) Model performance comparison showing faster convergence and lower final loss for multi-chip ensemble approaches. (b) Generalization error demonstrating reduced overfitting in multi-chip architectures. (c) Quantum error measurements under realistic noise conditions, showing the superior noise resilience of multi-chip ensembles compared to both standard single-chip VQCs and those with zero-noise extrapolation (ZNE). These results validate our theoretical predictions that controlled entanglement improves multiple aspects of QML performance simultaneously.
@@ 5.3 Trainability
Using the entangling capability measure [46], we quantified the quantum entanglement levels in single-chip and multi-chip ensemble VQCs with classical dimension reduction. For a total of 8 qubits, increasing the number of chips (i.e., dividing a large VQC into smaller subcircuits) reduced the overall entanglement while increasing the variance of gradients (Table 1).
Since barren plateaus are characterized by exponentially small gradient variance [13, 14], the higher gradient variance observed in multi-chip ensemble VQCs suggests a reduced risk of barren plateaus. These results, consistent with our theoretical analysis in Appendix C, demonstrate that the multi-chip ensemble approach can mitigate the risk of barren plateaus, thereby enhancing trainability.
@@ 6 Discussion
The multi-chip ensemble VQC framework addresses fundamental limitations of QML through distributed computation. Our experiments demonstrate simultaneous improvements in accuracy, generalization, noise resilience, and the ability to process high-dimensional data without dimension reduction.
Our approach provides a systematic architectural method for controlling entanglement, validating that balanced rather than maximal entanglement optimizes quantum advantage—aligning with recent work on barren plateaus and generalizability [50, 51, 52, 62]. The framework effectively navigates NISQ hardware constraints through distributed computational load across smaller circuits.
At a systems level, this method parallels data-parallel ensembles and dropout in deep learning: trading a monolithic model for robust, low-depth learners whose outputs are aggregated by a general function $g$. This creates a natural bridge between classical ensemble theory and quantum variational design, making our approach accessible to researchers beyond the QML community.
Limitations include our use of identical subcircuits across chips and non-optimized aggregation functions. Future work should explore heterogeneous subcircuit designs, task-specific aggregation strategies, and theoretical boundaries of quantum advantage preservation.
Our approach bridges theory with practical implementation, offering a pathway toward scalable QML that operates within current technological constraints while delivering advantages for complex learning tasks.